OpenStudy (anonymous):

which of the following is NOT equivalent to 45/70? 9/14 27/42 63/98 90/139

OpenStudy (callisto):

Simplify all the numbers given 45/70 = (9x5) /(5x14) = 9/14 Can you do the rest ?

OpenStudy (anonymous):

thanks alot this helped...

OpenStudy (callisto):

You're welcome :)